# Seeing What Matters: Visual Preference Policy Optimization for Visual Generation

### v2

@@ 1 Introduction
Reinforcement learning (RL) has recently emerged as an effective framework for aligning visual generative models [diffusion_ho, diffusion_song, flow_matching, rectifiedflow, liang2025integrating, zhang2024vast] with human preferences [firstrlddpm, fan2023dpok], enabling scalable supervision beyond paired data. Among RL-based approaches, Group Relative Policy Optimization (GRPO) [deepseek] has attracted attention for its group-wise comparison-based advantage formulation, which improves optimization stability and sample quality. Recent studies [dancegrpo, zhou2024flowgrpo] have successfully extended GRPO to diffusion and flow-based generators, confirming its potential for reinforcement-driven alignment in visual generation.
@@ 2 Related Work
RL for Visual Generation. Inspired by Proximal Policy Optimization (PPO) [ppo], early works [firstrlddpm, ddpo, fan2023dpok] integrated RL into diffusion models by optimizing the score function [scorebased] through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Recently, GRPO-based approaches [dancegrpo, zhou2024flowgrpo, mixgrpo, tempflow] have pushed visual generation to new heights. In particular, DanceGRPO [dancegrpo] and FlowGRPO [zhou2024flowgrpo] adapt GRPO to visual generation by reformulating Flow Matching’s [flow_matching] ODE sampling into an SDE formulation, enabling online RL training on state-of-the-art visual generative models. To further improve efficiency, MixGRPO [mixgrpo] introduces a mixed ODE-SDE strategy with a sliding window mechanism, significantly reducing training overhead while maintaining performance. However, all these methods overlook the inherent characteristics of visual content, which, unlike language, possesses rich spatial dimensions that could be exploited for more fine-grained optimization.
@@
Visual Perception Modeling. Modeling human visual perception has been a central theme in computer vision, with early approaches drawing direct inspiration from vision science. Saliency-based models [itti2001computational, jiang2015salicon] operationalized the idea that the visual system reduces scene complexity by prioritizing salient regions. Subsequent work [henderson2017meaning] highlighted the role of high-level semantics in guiding attention, leading to the notion of meaning maps, while eye-tracking studies [deepsaliency, henderson2018meaning] further revealed the non-uniform and dynamic nature of human gaze behavior. These perceptual insights have progressively shaped computational modeling, from the introduction of attention mechanisms in deep networks [visualattention], to perceptual loss [perceptualloss] which explicitly measures discrepancies between CNN feature maps to approximate human perceptual similarity, and more recently to robotics [emulating], where the adaptability of human vision inspired the Adaptive Vision Policy enabling agents to actively select optimal viewpoints.Visual preferences fundamentally rely on perceptual modeling. Building on this trajectory, we incorporate perceptual structuring into modern reinforcement learning for visual preference alignment, enabling content-adaptive optimization of visual content.
@@
Reward Model in Vision. A key bottleneck in applying RL to visual generation lies in the development of visual reward models. For image generation, recent works [pickscore, hpsv2, xu2023imagereward] have introduced perference-based reward models such as such as PickScore [pickscore], HPSv2 [hpsv2], and ImageReward [xu2023imagereward], which learn to predict human visual preferences. For video generation, VideoScore [he2024videoscore] introduces learnable metrics for direct evaluation, while VideoAlign [videoalign] assesses videos along three dimensions: visual quality, motion quality, and text alignment. More recently, VisionReward [visionreward] has been proposed as a fine-grained reward model for broader visual tasks. However, existing reward models primarily output scalar-level scores, which provide no information about where or why an image or video receives a high or low reward. More importantly, even though these models can capture fine-grained cues, a scalar reward collapses all spatial evidence into a single value. As a result, current GRPO-style alignment frameworks cannot exploit the rich spatial structure encoded in modern visual reward models.
@@ 3.1 Preliminaries
GRPO for Visual Generation. The denoising process of the diffusion and rectified flow can be formulated as a Markov Decision Process (MDP). Thus GRPO [deepseek] can be applied as following. Given a prompt $c$, the generative policy will sample a group of outputs $\{o_{1},o_{2},...,o_{G}\}$ with a group size of $G$ and optimize the policy model by maximizing the following objective function:
@@
SDE Sampling. State-of-the-art visual generative models increasingly adopt flow matching due to its efficiency and flexibility. However, flow matching typically relies on deterministic sampling based on an ordinary differential equation (ODE). The forward process in rectified flow [rectifiedflow] is defined as: $\mathrm{d}\mathbf{z}_{t}=\mathbf{u}_{t}\mathrm{d}t$, where $\mathbf{u}_{t}$ is the learned velocity field. The generative process reverses the ODE in time. However, GRPO requires stochastic exploration across multiple trajectory samples. To support RL within flow-matching frameworks, it becomes necessary to convert the ODE formulation into a stochastic differential equation (SDE).
@@ 3.2 Perceptual Structuring Module
Human visual preference is inherently selective and spatially biased [desimone1995neural, itti2001computational, henderson2017meaning]: observers focus on semantically informative areas while discounting redundant background. To capture this property, ViPO introduces a Perceptual Structuring Module (PSM) that extracts visual preference cues and encodes them into a preference allocation map used for structured advantage assignment. The PSM comprises a Visual Preference Extractor (VPE) and a Visual Preference Allocator (VPA).
@@ 4.1 Settings
Dataset. For image generation, we use the prompts from HPD [hpd]. The test set consists of 3200 prompts, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”. For video generation, we use the prompts from VidProM [vidprom] and randomly choose 1000 prompts as the test set, since VidProM does not provide a publicly released test split.
Backbones and Rewards. For image generation, we fine‑tune FLUX.1‑dev [blackforest2024flux] using HPSv2.1 [hpsv2] as the reward model, and further assess out‑of‑domain (OOD) generalization with PickScore [pickscore] and ImageReward [xu2023imagereward]. For video generation, we fine‑tune Wan2.1‑T2V‑14B‑480P [wan2.1] with VideoAlign [videoalign], which provides in‑domain reward signals for visual quality (VQ) and motion quality (MQ). OOD generalization is additionally evaluated on VBench [huang2024vbench].
@@ 4.2 Human Preference Reward
Quantitative Results. To validate the effectiveness of the proposed Visual Preference Policy Optimization (ViPO) in both image and video generation, we conduct comprehensive quantitative and qualitative experiments under human preference–based reward models. As DanceGRPO [dancegrpo] represents one of the most recent and widely adopted GRPO-based methods for visual generation with diffusion and flow-matching models, we adopt it as the baseline to provide a rigorous and representative evaluation. In addition, we further examine the impact of different visual backbones within the PSM.
The quantitative results of image generation are shown in Table 1. To assess the backbone sensitivity of ViPO, we construct three variants based on DINOv2 [oquab2023dinov2], SAM [sam], and ResNet [resnet], and all variants consistently outperform DanceGRPO across key metrics. Specifically, when HPS-v2.1 is used solely as the training reward model, ViPO achieves significant performance gains in both in-domain and out-domain evaluations. Among the variants, Among the variants, the DINO-based version performs the best, achieving the highest values in the in-domain HPSv2.1 and out-of-domain ImageReward. The ResNet-based variant exhibits unexpectedly good performance, particularly reaching the optimal value in the out-of-domain PickScore. The SAM-based variant is relatively weaker, but its metrics still surpass those of DanceGRPO. This may due to the features extracted by SAM being more inclined to low-level content rather than the high-level semantic information.
@@ A Details of PSM
Computation Details. The Perceptual Structuring Module (PSM) enriches the scalar reward signal by modeling human visual preference through pretrained vision backbones. As detailed in Section 3.2, PSM extracts perceptual features and constructs allocation maps that guide the redistribution of advantages across spatio-temporal locations. In the experiments, DINOv2 [oquab2023dinov2], ResNet [resnet], and SAM [sam] are respectively adopted as perceptual backbones for evaluating the effectiveness of PSM.
@@
For CNN-based models [resnet], intermediate feature maps $\mathbf{F}\in\mathbb{R}^{C\times H_{f}\times W_{f}}$ are extracted from a designated layer, where $\mathbf{F}_{c}\in\mathbb{R}^{H_{f}\times W_{f}}$ denotes the $c$-th channel. An activation map $\mathbf{S}\in\mathbb{R}^{H_{f}\times W_{f}}$ is obtained via channel-weighted aggregation:

### v3

@@ 1 Introduction
Reinforcement learning (RL) has recently emerged as an effective framework for aligning visual generative models [12, 27, 22, 24, 21, 38] with human preferences [3, 4], enabling scalable supervision beyond paired data. Among RL-based approaches, Group Relative Policy Optimization (GRPO) [5] has attracted attention for its group-wise comparison-based advantage formulation, which improves optimization stability and sample quality. Recent studies [36, 39] have successfully extended GRPO to diffusion and flow-based generators, confirming its potential for reinforcement-driven alignment in visual generation.
@@ 2 Related Work
RL for Visual Generation. Inspired by Proximal Policy Optimization (PPO) [26], early works [3, 1, 4] integrated RL into diffusion models by optimizing the score function [28] through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Recently, GRPO-based approaches [36, 39, 20, 8] have pushed visual generation to new heights. In particular, DanceGRPO [36] and FlowGRPO [39] adapt GRPO to visual generation by reformulating Flow Matching’s [22] ODE sampling into an SDE formulation, enabling online RL training on state-of-the-art visual generative models. To further improve efficiency, MixGRPO [20] introduces a mixed ODE-SDE strategy with a sliding window mechanism, significantly reducing training overhead while maintaining performance. However, all these methods overlook the inherent characteristics of visual content, which, unlike language, possesses rich spatial dimensions that could be exploited for more fine-grained optimization.
@@
Visual Perception Modeling. Modeling human visual perception has been a central theme in computer vision, with early approaches drawing direct inspiration from vision science. Saliency-based models [14, 15] operationalized the idea that the visual system reduces scene complexity by prioritizing salient regions. Subsequent work [10] highlighted the role of high-level semantics in guiding attention, leading to the notion of meaning maps, while eye-tracking studies [6, 11] further revealed the non-uniform and dynamic nature of human gaze behavior. These perceptual insights have progressively shaped computational modeling, from the introduction of attention mechanisms in deep networks [35], to perceptual loss [16] which explicitly measures discrepancies between CNN feature maps to approximate human perceptual similarity, and more recently to robotics [30], where the adaptability of human vision inspired the Adaptive Vision Policy enabling agents to actively select optimal viewpoints.Visual preferences fundamentally rely on perceptual modeling. Building on this trajectory, we incorporate perceptual structuring into modern reinforcement learning for visual preference alignment, enabling content-adaptive optimization of visual content.
@@
Reward Model in Vision. A key bottleneck in applying RL to visual generation lies in the development of visual reward models. For image generation, recent works [18, 32, 34] have introduced preference-based reward models such as such as PickScore [18], HPSv2 [32], and ImageReward [34], which learn to predict human visual preferences. For video generation, VideoScore [9] introduces learnable metrics for direct evaluation, while VideoAlign [23] assesses videos along three dimensions: visual quality, motion quality, and text alignment. More recently, VisionReward [33] has been proposed as a fine-grained reward model for broader visual tasks. However, existing reward models primarily output scalar-level scores, which provide no information about where or why an image or video receives a high or low reward. More importantly, even though these models can capture fine-grained cues, a scalar reward collapses all spatial evidence into a single value. As a result, current GRPO-style alignment frameworks cannot exploit the rich spatial structure encoded in modern visual reward models.
@@ 3.1 Preliminaries
GRPO for Visual Generation. The denoising process of the diffusion and rectified flow can be formulated as a Markov Decision Process (MDP). Thus GRPO [5] can be applied as following. Given a prompt $c$, the generative policy will sample a group of outputs $\{o_{1},o_{2},...,o_{G}\}$ with a group size of $G$ and optimize the policy model by maximizing the following objective function:
@@
SDE Sampling. State-of-the-art visual generative models increasingly adopt flow matching due to its efficiency and flexibility. However, flow matching typically relies on deterministic sampling based on an ordinary differential equation (ODE). The forward process in rectified flow [24] is defined as: $\mathrm{d}\mathbf{z}_{t}=\mathbf{u}_{t}\mathrm{d}t$, where $\mathbf{u}_{t}$ is the learned velocity field. The generative process reverses the ODE in time. However, GRPO requires stochastic exploration across multiple trajectory samples. To support RL within flow-matching frameworks, it becomes necessary to convert the ODE formulation into a stochastic differential equation (SDE).
@@ 3.2 Perceptual Structuring Module
Human visual preference is inherently selective and spatially biased [2, 14, 10]: observers focus on semantically informative areas while discounting redundant background. To capture this property, ViPO introduces a Perceptual Structuring Module (PSM) that extracts visual preference cues and encodes them into a preference allocation map used for structured advantage assignment. The PSM comprises a Visual Preference Extractor (VPE) and a Visual Preference Allocator (VPA).
@@ 4.1 Settings
Dataset. For image generation, we use the prompts from HPD [31]. The test set consists of 3200 prompts, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”. For video generation, we use the prompts from VidProM [37] and randomly choose 1000 prompts as the test set, since VidProM does not provide a publicly released test split.
Backbones and Rewards. For image generation, we fine‑tune FLUX.1‑dev [19] using HPSv2.1 [32] as the reward model, and further assess out‑of‑domain (OOD) generalization with PickScore [18] and ImageReward [34]. For video generation, we fine‑tune Wan2.1‑T2V‑14B‑480P [29] with VideoAlign [23], which provides in‑domain reward signals for visual quality (VQ) and motion quality (MQ). OOD generalization is additionally evaluated on VBench [13].
@@ 4.2 Human Preference Reward
Quantitative Results. To validate the effectiveness of the proposed Visual Preference Policy Optimization (ViPO) in both image and video generation, we conduct comprehensive quantitative and qualitative experiments under human preference–based reward models. As DanceGRPO [36] represents one of the most recent and widely adopted GRPO-based methods for visual generation with diffusion and flow-matching models, we adopt it as the baseline to provide a rigorous and representative evaluation. In addition, we further examine the impact of different visual backbones within the PSM.
The quantitative results of image generation are shown in Table 1. To assess the backbone sensitivity of ViPO, we construct three variants based on DINOv2 [25], SAM [17], and ResNet [7], and all variants consistently outperform DanceGRPO across key metrics. Specifically, when HPS-v2.1 is used solely as the training reward model, ViPO achieves significant performance gains in both in-domain and out-domain evaluations. Among the variants, the DINO-based version performs the best, achieving the highest values in the in-domain HPSv2.1 and out-of-domain ImageReward. The ResNet-based variant exhibits unexpectedly good performance, particularly reaching the optimal value in the out-of-domain PickScore. The SAM-based variant is relatively weaker, but its metrics still surpass those of DanceGRPO. This may due to the features extracted by SAM being more inclined to low-level content rather than the high-level semantic information.
@@ A Details of PSM
Computation Details. The Perceptual Structuring Module (PSM) enriches the scalar reward signal by modeling human visual preference through pretrained vision backbones. As detailed in Section 3.2, PSM extracts perceptual features and constructs allocation maps that guide the redistribution of advantages across spatio-temporal locations. In the experiments, DINOv2 [25], ResNet [7], and SAM [17] are respectively adopted as perceptual backbones for evaluating the effectiveness of PSM.
@@
For CNN-based models [7], intermediate feature maps $\mathbf{F}\in\mathbb{R}^{C\times H_{f}\times W_{f}}$ are extracted from a designated layer, where $\mathbf{F}_{c}\in\mathbb{R}^{H_{f}\times W_{f}}$ denotes the $c$-th channel. An activation map $\mathbf{S}\in\mathbb{R}^{H_{f}\times W_{f}}$ is obtained via channel-weighted aggregation:
@@
Training Overhead. To further quantify the computational overhead introduced by PSM, we report the training overhead of ViPO with different perceptual backbones in Table 5. PSM introduces only minor overhead relative to the original GRPO training pipeline: the training step time increases by only 1.0–1.8% for image generation and 4.8% for video generation, while the increase in peak memory usage remains below 4.5%. The detailed cost analysis (forward / PCA / resample) indicates that PSM incurs negligible latency relative to the total training step, without modifying the generative backbone.
Table 5: Training overhead of ViPO with different PSM backbones, compared with the corresponding DanceGRPO baseline.
Method Step Time (s) PSM/sample (ms) Peak Mem (MB)
Flux+DanceGRPO 59.66 47084
Flux+ViPO (DINO) 60.47 (+1.4%) 49.73 (27.9/21.4/0.5) 48964 (+4.0%)
Flux+ViPO (ResNet) 60.24 (+1.0%) 4.22 (3.7/0.2/0.4) 47200 (+0.3%)
Flux+ViPO (SAM) 60.75 (+1.8%) 33.25 (20.5/12.2/0.6) 47422 (+0.7%)
Wan+DanceGRPO 203.45 25498
Wan+ViPO (DINO) 213.27 (+4.8%) 972.5 (413/498/17.5) 26654 (+4.5%)